# Effect of Synthetic Jets Actuator Parameters on Deep Reinforcement Learning-Based Flow Control Performance in a Square Cylinder

### v3

@@ I INTRODUCTION
In the backdrop of advancements in sophisticated algorithms, computational hardware, open-source software and the integration of vast amounts of data in the era of big data, scientific inquiry is transitioning from first principles to data-driven methods.DataDriven; Datadrivenscience The field of fluid dynamics is also benefiting from this trend, with machine learning offering sophisticated algorithms capable of handling high-dimensional, large-scale fluid dynamics data,Brunton, Noack, and Koumoutsakos (2020) establishing mathematical modeling frameworks, and providing integrated, modular packages and libraries as open-source resources.Brunton2015; Stevendynamical Machine learning excels in tasks like feature extraction, modal analysis, and data mining, providing techniques for extracting insights from datasets.Batta (2020); Datadrivenscience Experimental fluid dynamics measurements and numerical simulation generate extensive datasets, making them ideal candidates for techniques capable of handling high-dimensional, large-scale data.SCOTTCOLLIS2004237; DataDriven Furthermore, machine learning algorithms can embed physical information from the field of fluid dynamics and adaptively perform tasks related to execution, control, and optimization.Multifidelity; Bewley2001 The field of flow control is rapidly advancing, propelled by the development of state-of-the-art machine learning algorithms and unprecedented volumes of data from high-precision numerical simulations across multiple spatiotemporal scales.Brunton, Noack, and Koumoutsakos (2020); Multifidelity; Bewley2001
@@
Active flow control (AFC) is an advanced fluid management strategy aimed at enhancing fluid properties and performance through proactive interventions in the flow field. By introducing energy or momentum, AFC alters the fluid’s natural state, employing techniques such as fluid injection and suction, synthetic jets, and electromagnetic control.annurevCattafesta; Aram2018 Synthetic jet technology produces jets by cyclically inhaling and expelling fluid from a fixed location, requiring no external energy source.annurevfluidGlezer; SCOTTCOLLIS2004237 This method offers several advantages, including independence from external fluid sources, high controllability, low energy consumption, and suitability for operation in narrow or complex spaces.jahanmiri2010active; Smith1998 These advantages render synthetic jets particularly promising tools for present and future flow control technologies.MAL071; Aerospace2023 AFC is widely applied across various sectors, including aerospace, automotive, energy production, and environmental engineering, where it serves to reduce drag, control vortices, enhance lift, and stabilize flows.SCOTTCOLLIS2004237; smith2003comparison Despite its considerable advantages, the implementation and practical application of AFC face challenges due to the system’s strong nonlinearity, high dimensionality, and time delays. Recent advancements in machine learning algorithms and control theory are propelling progress in adaptive, real-time intelligent control, facilitating more effective management of complex flow dynamics and further augmenting the capabilities of AFC technologies.Brunton2015; Datadrivenscience; Stevendynamical
@@
Deep Reinforcement Learning (DRL) merges the capabilities of Deep Learning (DL) for processing high-dimensional data with the decision-making and policy optimization strengths of Reinforcement Learning (RL).mnih2013playing; lecun2015deep DRL is recognized for its ability to learn and adapt autonomously, enabling end-to-end learning that is highly applicable across complex environments. DRL has demonstrated exceptional performance in a wide range of domains, including game AI, robotic control, autonomous driving, resource management and scheduling, and financial trading. It is capable of autonomously learning optimal decision-making strategies in complex environments, surpassing human-level capabilities.mnih2015human; lillicrap2019continuous; henderson2019deep DRL is a sophisticated blend of DL and RL, two powerful branches of machine learning. It combines the feature extraction capabilities of DL for complex high-dimensional data with the decision-making abilities of RL through interactions with the environment.li2018deep; kaiser2024modelbased This synergy allows DRL to excel in environments that require both advanced perception and precise control. Despite challenges related to sample efficiency and stability, DRL demonstrates remarkable potential on a variety of challenging decision-making problems.Arulkumaran et al. (2017); kaiser2024modelbased
@@
Given the significant potential of DRL in control and optimization, a large number of scholars have been applying it in the field of flow control. rabault2019artificialrabault2019artificial pioneered the application of DRL to AFC technologies, achieving an 8% reduction in drag around a cylinder at $Re=100$. This seminal work sparked interest in applying DRL techniques within the field of fluid dynamics. tang2020robusttang2020robust focused on the robustness across different Reynolds numbers, while parisRobustFlowControl2021parisRobustFlowControl2021 explored the impact of probe distribution on control performance. Li and ZhangLi and Zhang (2022) applied DRL-based AFC to cylinders under various blockage ratios, integrating flow physics mechanisms during the control process to achieve significant outcomes. DixiapnasDixiapnas developed an active flow control strategy based on DRL through experimental fluid dynamics, demonstrating the effectiveness of applying reinforcement learning in experimental fluid mechanics. rabault2019acceleratingrabault2019accelerating and Jia and XuJia and Xu (2024) addressed issues related to parallel strategies and training when coupling Computational Fluid Dynamics (CFD) codes with DRL framework in parallel computing environments. Wang et al.Wang et al. (2022) developed an open source library that integrated CFD code OpenFoam with DRL framework TensorForce. DRL demonstrates excellent control performance in the field of AFC, sparking increasing exploration. There are applications from low to relatively high Reynolds numbers, transitions from 2D to 3D configurations, shifts from numerical simulations to fluid experiments, and geometric changes in the body of interest from circular to square bodies. While there is a plethora of related work, we will not delve into each individually. Instead, we summarize in Tab. 1 some relevant research on the application of DRL-based AFC methods in the context of flow around bluff bodies.
@@
100 Cylinder rabault2019artificialrabault2019artificial Synthetic Jets PPO FeniCS 8% -
@@
100 Cylinder castellanoscastellanos Synthetic Jets LGPC/PPO FeniCS 8% -
100 Cylinder jia2024deepjia2024deep Synthetic Jets PPO OpenFOAM 8% YES
100 Cylinder ren2021bluffren2021bluff WSLB PPO LBM - -
120 Cylinder parisRobustFlowControl2021parisRobustFlowControl2021 Synthetic Jets S-PPO-CMA FastS 18.4% -
@@
100-300 Cylinder he2023policyhe2023policy Synthetic Jets PPO OpenFOAM 6%-24% -
100-400 Cylinder tangRobusttangRobust Synthetic Jets PPO FeniCS 6%-39% -
@@
10,160 Cylinder DixiapnasDixiapnas Cylinders TD3 Lilypad 30% -
100 Ellipse jia2024deepjia2024deep Synthetic Jets PPO OpenFOAM 16% YES
@@
100 Square Xia2024Xia2024 Synthetic Jets SAC and TQC FEniCS 17.3% -
100-400 Square jia2024robustjia2024robust Synthetic Jets SAC OpenFOAM 14%-47% YES
@@
The flow around a square cylinder exhibits instabilities similar to those around a circular cylinder, but with different separation mechanisms and variations in lift, drag, and Strouhal number with Reynolds number. For square cylinders, the separation points are fixed at the leading or trailing edge, depending on the Reynolds number.Sohankar, Norberg, and Davidson (1998) The location of these separation points is crucial for guiding active flow control strategies. Synthetic jets control flow separation by periodically injecting or extracting fluid into or from the boundary layer, directly altering its momentum distribution.annurevCattafesta; YOU20081349 This enhances the energy within the boundary layer, delays the separation point, and prevents flow separation. The effectiveness of synthetic jets depends largely on parameters such as amplitude, frequency, and application location.SCOTTCOLLIS2004237; Aram2018 Extensive parameter studies are necessary to optimize control performance. The location of the synthetic jet is particularly crucial. Selecting the appropriate location maximizes the jet’s impact on the boundary layer, delays the separation point, and improves control performance. By optimizing the amplitude and frequency of the jet and selecting the best application location, optimal flow control can be achieved, significantly enhancing aerodynamic performance and system stability.
@@ II.1 Flow environment
In the current study, the incompressible flow dynamics are analyzed using the open-source CFD toolkit, OpenFOAM, as outlined by Jasak et al. jasakLibrary; jasak2009 OpenFOAM stands as a rigorously tested and extensively applied computational framework that offers a comprehensive suite of numerical algorithms for the resolution of the Navier-Stokes equations. The solver implemented within OpenFOAM employs the finite volume method for the discretization of the computational domain, subdividing it into a structured mesh of control volumes. The numerical simulations use the pimpleFoam solver to solve the unsteady Navier-Stokes equations. The temporal discretization employs the Backward scheme, while the divergence discretization of the velocity field uses the Gauss linear scheme, and the gradient discretization of the velocity utilizes the leastSquares method. These discretization schemes are selected to ensure the requisite mathematical precision and stability for the simulation of fluid flow dynamics. To promote numerical stability within the simulations, the time step size is meticulously selected as $\Delta t=0.0005$.
@@
sen2011flowsen2011flow - 1.530 - - 0.145
SharmaSharma - 1.490 - - 0.150
SinghSingh - 1.510 - - 0.147
@@
ShearShear - 2.050 - - 0.133
DependenceDependence - 1.940 - - 0.136
@@ II.2 Machine learning
The agent interacts with the environment, selecting actions according to its current policy $\pi$. The experience tuples $(s_{i},a_{i},r_{i},s_{i+1})$ collected from these interactions are stored in a replay buffer. When updating the network parameters, the algorithm samples batches of experience data from the replay buffer. Moreover, entropy ($\mathcal{H}$) is a metric that quantifies the uncertainty of a random variable. Within the framework of reinforcement learning, the entropy of a policy, $\mathcal{H}(\pi(\cdot\mid s_{i}))$, represents the stochasticity of the policy at a given state $s_{i}$.haarnoja2018soft; Haarnoja et al. (2018, 2019a) The expression for this is given by:
@@
where $\mathcal{H}(\cdot)$ is the entropy measure and the variable $\alpha$ functions as a temperature parameter that calibrates the relative significance of the entropy term against the reward, thereby modulating the stochasticity of the resulting optimal policy.Haarnoja et al. (2018, 2019b)
@@ II.3 DRL-Enhanced Active Flow Control
where $C_{D,0}$ represents the baseline drag coefficient $C_{D}$, serving as a reference point. $(C_{D})_{T_{i}}$ denotes the drag coefficient at time step $T_{i}$, with the objective being to minimize this value relative to the baseline. $(C_{L})_{T_{i}}$ denotes the lift coefficient at time step $T_{i}$. The parameter $\omega$ is a weight factor for the lift coefficient. We adopt a penalty factor of $0.2$ based on rabault2019artificialrabault2019artificial, balancing drag reduction and lift suppression.
@@ III.1 Baseline flow simulation
In Fig. 9, simulations are performed for the cases of $Re=100$ and $Re=500$, where the wake flow behind the semi-square cylinder is in a steady state without vortex shedding, and the only reason for pressure drag is flow separation. The total drag coefficient encapsulates the intrinsic drag associated with the baseline flow and the correctable component, with only the second part of drag being adjustable through strategic manipulation of the von Kármán vortex sheddingprotas2002drag. We compute the asymptotic drag coefficients, denoted as $C_{Dh}$, for the semi-square cylinder at $Re=100$ and $Re=500$, which are found to be 0.668 and 0.498, respectively. These values are extended to obtain the baseline drag coefficients, denoted as $C_{Db}$, for a fully square cylinder in a steady-state flow, resulting in values of 1.335 and 0.995 for $Re=100$ and $Re=500$, respectively. The result of $Re=100$ is consistent with the study of Xia2024Xia2024. The baseline drag coefficient, $C_{Db}$, is compared with the drag obtained during active flow control to evaluate the effectiveness of the control strategy in reducing drag.
@@ III.3.1 Location of Synthetic Jets
Our research indicates that for the baseline flow at $Re=100$, positioning the synthetic jets near the trailing corner yields a maximum drag reduction of 14.4%. For the baseline flow at $Re=500$, the optimal drag reduction efficiency is achieved by placing the synthetic jet near the leading corner, with a maximum drag reduction rate of 65.5%. This phenomenon is related to the differing locations of flow separation around the square cylinder at $Re=100$ and $Re=500$. At $Re=100$, flow separation occurs near the trailing corner, whereas at $Re=500$, flow separation occurs near the leading corner, with the separated bubble reattaching to the sidewall and subsequently reattaching near the trailing corner. Based on the flow control effects given by DRL, it is found that placing the synthetic jets at the location of flow separation is more conducive to achieving control objectives. Similarly, KIM2009172 and KIM2009172 observed that when controlling flow separation on an airfoil using synthetic jets, the control effectiveness is optimal when the separation point coincides with the location of the synthetic jets.

### v4

@@ I INTRODUCTION
In the backdrop of advancements in sophisticated algorithms, computational hardware, open-source software and the integration of vast amounts of data in the era of big data, scientific inquiry is transitioning from first principles to data-driven methods.Manohar et al. (2018); Luchtenburg (2021) The field of fluid dynamics is also benefiting from this trend, with machine learning offering sophisticated algorithms capable of handling high-dimensional, large-scale fluid dynamics data,Brunton, Noack, and Koumoutsakos (2020) establishing mathematical modeling frameworks, and providing integrated, modular packages and libraries as open-source resources.Brunton and Noack (2015); Brunton, Proctor, and Kutz (2016) Machine learning excels in tasks like feature extraction, modal analysis, and data mining, providing techniques for extracting insights from datasets.Mahesh (2020); Luchtenburg (2021) Experimental fluid dynamics measurements and numerical simulation generate extensive datasets, making them ideal candidates for techniques capable of handling high-dimensional, large-scale data.Scott Collis et al. (2004); Manohar et al. (2018) Furthermore, machine learning algorithms can embed physical information from the field of fluid dynamics and adaptively perform tasks related to execution, control, and optimization.Perdikaris, Venturi, and Karniadakis (2016); Bewley, Moin, and Temam (2001) The field of flow control is rapidly advancing, propelled by the development of state-of-the-art machine learning algorithms and unprecedented volumes of data from high-precision numerical simulations across multiple spatiotemporal scales.Brunton, Noack, and Koumoutsakos (2020); Perdikaris, Venturi, and Karniadakis (2016); Bewley, Moin, and Temam (2001)
@@
Active flow control (AFC) is an advanced fluid management strategy aimed at enhancing fluid properties and performance through proactive interventions in the flow field. By introducing energy or momentum, AFC alters the fluid’s natural state, employing techniques such as fluid injection and suction, synthetic jets, and electromagnetic control.Cattafesta and Sheplak (2011); Aram et al. (2018) Synthetic jet technology produces jets by cyclically inhaling and expelling fluid from a fixed location, requiring no external energy source.Glezer and Amitay (2002); Scott Collis et al. (2004) This method offers several advantages, including independence from external fluid sources, high controllability, low energy consumption, and suitability for operation in narrow or complex spaces.Jahanmiri (2010); Smith and Glezer (1998) These advantages render synthetic jets particularly promising tools for present and future flow control technologies.François-Lavet et al. (2018); Xie et al. (2023) AFC is widely applied across various sectors, including aerospace, automotive, energy production, and environmental engineering, where it serves to reduce drag, control vortices, enhance lift, and stabilize flows.Scott Collis et al. (2004); Smith and Swift (2003) Despite its considerable advantages, the implementation and practical application of AFC face challenges due to the system’s strong nonlinearity, high dimensionality, and time delays. Recent advancements in machine learning algorithms and control theory are propelling progress in adaptive, real-time intelligent control, facilitating more effective management of complex flow dynamics and further augmenting the capabilities of AFC technologies.Brunton and Noack (2015); Luchtenburg (2021); Brunton, Proctor, and Kutz (2016)
@@
Deep Reinforcement Learning (DRL) merges the capabilities of Deep Learning (DL) for processing high-dimensional data with the decision-making and policy optimization strengths of Reinforcement Learning (RL).Mnih et al. (2013); LeCun, Bengio, and Hinton (2015) DRL is recognized for its ability to learn and adapt autonomously, enabling end-to-end learning that is highly applicable across complex environments. DRL has demonstrated exceptional performance in a wide range of domains, including game AI, robotic control, autonomous driving, resource management and scheduling, and financial trading. It is capable of autonomously learning optimal decision-making strategies in complex environments, surpassing human-level capabilities.Mnih et al. (2015); Lillicrap et al. (2019); Henderson et al. (2019) DRL is a sophisticated blend of DL and RL, two powerful branches of machine learning. It combines the feature extraction capabilities of DL for complex high-dimensional data with the decision-making abilities of RL through interactions with the environment.Li (2018); Kaiser et al. (2024) This synergy allows DRL to excel in environments that require both advanced perception and precise control. Despite challenges related to sample efficiency and stability, DRL demonstrates remarkable potential on a variety of challenging decision-making problems.Arulkumaran et al. (2017); Kaiser et al. (2024)
@@
Given the significant potential of DRL in control and optimization, a large number of scholars have been applying it in the field of flow control. Rabault et al.Rabault et al. (2019) pioneered the application of DRL to AFC technologies, achieving an 8% reduction in drag around a cylinder at $Re=100$. This seminal work sparked interest in applying DRL techniques within the field of fluid dynamics. Tang et al.Tang et al. (2020a) focused on the robustness across different Reynolds numbers, while Paris, Beneddine, and DandoisParis, Beneddine, and Dandois (2021) explored the impact of probe distribution on control performance. Li and ZhangLi and Zhang (2022) applied DRL-based AFC to cylinders under various blockage ratios, integrating flow physics mechanisms during the control process to achieve significant outcomes. Fan et al.Fan et al. (2020) developed an active flow control strategy based on DRL through experimental fluid dynamics, demonstrating the effectiveness of applying reinforcement learning in experimental fluid mechanics. Rabault and KuhnleRabault and Kuhnle (2019a) and Wang and XuWang and Xu (2024a) addressed issues related to parallel strategies and training when coupling Computational Fluid Dynamics (CFD) codes with DRL framework in parallel computing environments. Wang et al.Wang et al. (2022) developed an open source library that integrated CFD code OpenFoam with DRL framework TensorForce. DRL demonstrates excellent control performance in the field of AFC, sparking increasing exploration. There are applications from low to relatively high Reynolds numbers, transitions from 2D to 3D configurations, shifts from numerical simulations to fluid experiments, and geometric changes in the body of interest from circular to square bodies. While there is a plethora of related work, we will not delve into each individually. Instead, we summarize in Tab. 1 some relevant research on the application of DRL-based AFC methods in the context of flow around bluff bodies.
@@
100 Cylinder Rabault et al.Rabault et al. (2019) Synthetic Jets PPO FeniCS 8% -
@@
100 Cylinder Castellanos et al.Castellanos et al. (2022) Synthetic Jets LGPC/PPO FeniCS 8% -
100 Cylinder Jia and XuJia and Xu (2024) Synthetic Jets PPO OpenFOAM 8% YES
100 Cylinder Ren, Wang, and TangRen, Wang, and Tang (2021) WSLB PPO LBM - -
120 Cylinder Paris, Beneddine, and DandoisParis, Beneddine, and Dandois (2021) Synthetic Jets S-PPO-CMA FastS 18.4% -
@@
100-300 Cylinder He et al.He et al. (2023) Synthetic Jets PPO OpenFOAM 6%-24% -
100-400 Cylinder Tang et al.Tang et al. (2020b) Synthetic Jets PPO FeniCS 6%-39% -
@@
10,160 Cylinder Fan et al.Fan et al. (2020) Cylinders TD3 Lilypad 30% -
100 Ellipse Jia and XuJia and Xu (2024) Synthetic Jets PPO OpenFOAM 16% YES
@@
100 Square Xia et al.Xia et al. (2024) Synthetic Jets SAC and TQC FEniCS 17.3% -
100-400 Square Wang and XuWang and Xu (2024b) Synthetic Jets SAC OpenFOAM 14%-47% YES
@@
The flow around a square cylinder exhibits instabilities similar to those around a circular cylinder, but with different separation mechanisms and variations in lift, drag, and Strouhal number with Reynolds number. For square cylinders, the separation points are fixed at the leading or trailing edge, depending on the Reynolds number.Sohankar, Norberg, and Davidson (1999) The location of these separation points is crucial for guiding active flow control strategies. Synthetic jets control flow separation by periodically injecting or extracting fluid into or from the boundary layer, directly altering its momentum distribution.Cattafesta and Sheplak (2011); You and Moin (2008) This enhances the energy within the boundary layer, delays the separation point, and prevents flow separation. The effectiveness of synthetic jets depends largely on parameters such as amplitude, frequency, and application location.Scott Collis et al. (2004); Aram et al. (2018) Extensive parameter studies are necessary to optimize control performance. The location of the synthetic jet is particularly crucial. Selecting the appropriate location maximizes the jet’s impact on the boundary layer, delays the separation point, and improves control performance. By optimizing the amplitude and frequency of the jet and selecting the best application location, optimal flow control can be achieved, significantly enhancing aerodynamic performance and system stability.
@@ II.1 Flow environment
In the current study, the incompressible flow dynamics are analyzed using the open-source CFD toolkit, OpenFOAM, as outlined by Jasak et al. Jasak et al. (2007); Jasak (2009) OpenFOAM stands as a rigorously tested and extensively applied computational framework that offers a comprehensive suite of numerical algorithms for the resolution of the Navier-Stokes equations. The solver implemented within OpenFOAM employs the finite volume method for the discretization of the computational domain, subdividing it into a structured mesh of control volumes. The numerical simulations use the pimpleFoam solver to solve the unsteady Navier-Stokes equations. The temporal discretization employs the Backward scheme, while the divergence discretization of the velocity field uses the Gauss linear scheme, and the gradient discretization of the velocity utilizes the leastSquares method. These discretization schemes are selected to ensure the requisite mathematical precision and stability for the simulation of fluid flow dynamics. To promote numerical stability within the simulations, the time step size is meticulously selected as $\Delta t=0.0005$.
@@
Sen, Mittal, and BiswasSen, Mittal, and Biswas (2011) - 1.530 - - 0.145
Sharma and EswaranSharma and Eswaran (2004) - 1.490 - - 0.150
Singh et al.Singh et al. (2009) - 1.510 - - 0.147
@@
Cao, Ge, and TamuraCao, Ge, and Tamura (2012) - 2.050 - - 0.133
Bai and AlamBai and Alam (2018) - 1.940 - - 0.136
@@ II.2 Machine learning
The agent interacts with the environment, selecting actions according to its current policy $\pi$. The experience tuples $(s_{i},a_{i},r_{i},s_{i+1})$ collected from these interactions are stored in a replay buffer. When updating the network parameters, the algorithm samples batches of experience data from the replay buffer. Moreover, entropy ($\mathcal{H}$) is a metric that quantifies the uncertainty of a random variable. Within the framework of reinforcement learning, the entropy of a policy, $\mathcal{H}(\pi(\cdot\mid s_{i}))$, represents the stochasticity of the policy at a given state $s_{i}$.Haarnoja et al. (2018a, b, 2019a) The expression for this is given by:
@@
where $\mathcal{H}(\cdot)$ is the entropy measure and the variable $\alpha$ functions as a temperature parameter that calibrates the relative significance of the entropy term against the reward, thereby modulating the stochasticity of the resulting optimal policy.Haarnoja et al. (2018b, 2019b)
@@ II.3 DRL-Enhanced Active Flow Control
where $C_{D,0}$ represents the baseline drag coefficient $C_{D}$, serving as a reference point. $(C_{D})_{T_{i}}$ denotes the drag coefficient at time step $T_{i}$, with the objective being to minimize this value relative to the baseline. $(C_{L})_{T_{i}}$ denotes the lift coefficient at time step $T_{i}$. The parameter $\omega$ is a weight factor for the lift coefficient. We adopt a penalty factor of $0.2$ based on Rabault et al.Rabault et al. (2019), balancing drag reduction and lift suppression.
@@ III.1 Baseline flow simulation
In Fig. 9, simulations are performed for the cases of $Re=100$ and $Re=500$, where the wake flow behind the semi-square cylinder is in a steady state without vortex shedding, and the only reason for pressure drag is flow separation. The total drag coefficient encapsulates the intrinsic drag associated with the baseline flow and the correctable component, with only the second part of drag being adjustable through strategic manipulation of the von Kármán vortex sheddingProtas and Wesfreid (2002). We compute the asymptotic drag coefficients, denoted as $C_{Dh}$, for the semi-square cylinder at $Re=100$ and $Re=500$, which are found to be 0.668 and 0.498, respectively. These values are extended to obtain the baseline drag coefficients, denoted as $C_{Db}$, for a fully square cylinder in a steady-state flow, resulting in values of 1.335 and 0.995 for $Re=100$ and $Re=500$, respectively. The result of $Re=100$ is consistent with the study of Xia et al.Xia et al. (2024). The baseline drag coefficient, $C_{Db}$, is compared with the drag obtained during active flow control to evaluate the effectiveness of the control strategy in reducing drag.
@@ III.3.1 Location of Synthetic Jets
Our research indicates that for the baseline flow at $Re=100$, positioning the synthetic jets near the trailing corner yields a maximum drag reduction of 14.4%. For the baseline flow at $Re=500$, the optimal drag reduction efficiency is achieved by placing the synthetic jet near the leading corner, with a maximum drag reduction rate of 65.5%. This phenomenon is related to the differing locations of flow separation around the square cylinder at $Re=100$ and $Re=500$. At $Re=100$, flow separation occurs near the trailing corner, whereas at $Re=500$, flow separation occurs near the leading corner, with the separated bubble reattaching to the sidewall and subsequently reattaching near the trailing corner. Based on the flow control effects given by DRL, it is found that placing the synthetic jets at the location of flow separation is more conducive to achieving control objectives. Similarly, Kim and Kim (2009) and Kim and Kim observed that when controlling flow separation on an airfoil using synthetic jets, the control effectiveness is optimal when the separation point coincides with the location of the synthetic jets.